A daredevil is shot out of a cannon at 45.0° to the horizontal with an initial speed of 29.5 m/s. A net is positioned a horizontal distance of 50.5 m from the cannon. At what height above the cannon should the net be placed in order to catch the daredevil?
1
Expert's answer
2019-10-11T10:47:46-0400

"h = x tg\alpha - \frac{gx^2}{2v^2 cos^2 \alpha}"

Where x -distance, alpha - angle to horizont, v - speed, h height where should be net placed


"h = 43.31 m"